OpenStudy (anonymous):

if a system has an infinite answer of solutions,use set builder notation to write the solution. if a system has no solution state this. 4p*2q=16 5p+7q=1

OpenStudy (anonymous):

Did you mean 4p+2q=16?

OpenStudy (anonymous):

$$4p+2q=16\\2p+q=8\\\ \\5p+7q=1$$If we multiply \(2p+q=8\) by \(-7\), we get \(-14p-7q=-56\). Combine that with our second equation, \(5p+7q=1\), and we eliminate the dependency on \(q\) to yield:$$-9p=-55\\p=\frac{55}9$$Now, we can substitute into our earlier equation, \(2p+q=8\), to solve for \(q\).$$\frac{110}9+q=8=\frac{72}9\\q=\frac{72-110}9=-\frac{38}{9}$$Thus our solutions \((p,q)\in\{(\frac{55}9,-\frac{38}9)\}\)